OpenStudy (anonymous):

evaluate the following expression for x=5 and y=-7 5x+7y-3x+8 A. -31 B.-16 C.-7 D.29

OpenStudy (anonymous):

5(5)+7(-7)-3(5)+8= 25-49-15+8=-31

OpenStudy (anonymous):

A